# DanceChat: Large Language Model-Guided Music-to-Dance Generation

### v1

@@ 4.3. Ablation Study
M B T PFC $\downarrow$ Divk $\rightarrow$ Beat Align. Score $\uparrow$
1.536 9.48 0.26
@@
Table 4. Ablation study of modalities, where M indicates music, B indicates beat and T indicates text. The best results are indicated in bold, and the second best are underlined.
5. Conclusion
In this paper, we present DanceChat, a multi-modal, language-guided framework for music-to-dance generation, addressing the significant semantic gap between music and dance. By leveraging a Large Language Model (LLM) as a pseudo-choreographer, our approach introduces natural language as an intermediary between music and motion, transforming abstract musical elements into concrete dance instructions. Our proposed model architecture comprises three core components: Pseudo Instruction Generation, which translates musical meta-information into text-based dance instructions; Multi-modal Fusion, which integrates music, beat, and textual guidance into a cohesive representation; and Diffusion-based Motion Synthesis, which employs a denoising process to generate smooth, expressive dance motions. Through comprehensive experiments on the AIST++ dataset, DanceChat demonstrates competitive performance across multiple metrics. The results confirm that each modality contributes uniquely to dance quality, while our multi-modal alignment loss plays a crucial role in harmonizing these modalities to produce musically aligned sequences.
Limitations and Future Work. While DanceChat achieves strong results, certain limitations remain. First, the reliance on pre-trained LLMs and external encoders introduces a dependency on existing models that may not be fully optimized for dance generation. Future work could explore custom training for language models specifically tailored for dance, potentially enhancing the semantic richness of generated instructions.

### v2

@@ 4.3 Ablation Study
M B T PFC $\downarrow$ Divk $\rightarrow$ BAS $\uparrow$
1.536 9.48 0.23
@@
Table 3: Ablation study of modalities, where M indicates music, B indicates beat and T indicates text. The best results are indicated in bold, and the second best are underlined.
5 Conclusion
We present DanceChat, a novel multi-modal framework that leverages language as an intermediate modality to bridge the semantic gap between music and dance. By incorporating textual dance instructions generated by LLMs and integrating them with both music and beat features, our approach enables semantically guided and rhythmically aligned motion generation. Comprehensive experiments on AIST++ dataset demonstrate that DanceChat achieves strong performance across multiple aspects of dance quality, including physical plausibility, motion diversity, and beat synchronization. Ablation results confirm that each modality contributes uniquely to dance quality, while our multi-modal alignment loss plays a crucial role in harmonizing these modalities to produce musically aligned motion sequences. Beyond quantitative improvements, DanceChat offers a new perspective on choreography modeling, where natural language serves not only as descriptive guidance but also as a bridge across modalities.